A box contains (π‘₯+1) red marbles, (2π‘₯+4) yellow marbles, and (4π‘₯βˆ’6) green marbles


(i)

Find an expression, in terms of π‘₯, for the total number of marbles in the box.


(ii)

A marble is drawn at random from the box. Write down an expression, in terms of π‘₯, for the probability that the marble is red.


(iii)

Given that probability in part (ii) is 1/5, find the value of π‘₯.



1
Expert's answer
2022-03-31T08:10:02-0400

(i)

The total number of marbles in the box:

"N=x+1+2x+4+4x-6=7x-1."


(ii)

The probability that the drawn marble is red:

"P=\cfrac{x+1}{7x-1}."


(iii)

"\cfrac{x+1}{7x-1}=\cfrac{1}{5},\\
5x+5=7x-1,\\
2x=6,\\
x=3."